Grade 12th passPhysical Chemistry

100cc of 1.5% solution of urea is found to have an osmotic pressure of 6atm and 100cc of 3.42% solution of cane sugar is found to have an osmotic pressure of 2.4atm. If two solutions are mixed, the osmotic pressure of the resulting mix is?
a)8.4 atm b)4.2 atm c)16.8 atm d)2.1atm

Ashmeet Lamba
9 Years agoGrade 12th pass

2 Answers

Vikas TU
ApprovedApproved Tutor Answer9 Years ago
caclulate the mlolarity first for both solution.
Urea = > (1.5*1000/M*100)
Cane Sugar => 3.42*1000/m*100
Where M and m are there molar masses respectively.
Thus net molarity would be from M1V1 = M2V2.
And net temp. would be T1 and T2 from both O.P. respectively.
Fianl O.P = 1*C*S*(T1 + T2).
C is net molarity of soltion after mixing and S is R constant)
Faraz Hashmi
6 Years ago
Pressure of urea= 6/2 = 3 atm
Pressure of canesugar= 2.4/2 =1.2 atm
We can see that both the reactions undergo at same temperature. Therefore its an ideal condition. 
Therefore, p[net]= 3+1.2 =  4.2 atm i.e. option {b}